# Particle production and hadronization temperature of the massive Schwinger model

### v1

@@ VII Conclusions
In this work we studied the dynamics and particle production from a string between a quark-antiquark pair produced, for example, in an electron-positron collision. As a model we used the bosonized version of the massive Schwinger model, which makes dynamics particularly easy to simulate numerically. We examine the dynamics of the coherent field and the particle production. Remarkably, the total particle number per rapidity interval to the fermion mass for highly massive or weakly coupled fermions turns out to be well represented by a Boltzmann factor.
@@
Very interesting extensions of this work include the consideration of flavor and color numbers, the incorporation of the production description beyond mesons, and taking into account baryons. Moreover, for the bounce scenario discussed in Sec. IV the precise determination of the quantum state of fluctuations after the bounce would be highly valuable. The question of whether this can be achieved through analytic continuation from a calculation in Euclidean space is left for future work.
@@ ACKNOWLEDGEMENTS
For discussions on related topics, the authors would like to thank Sebastian Erne, Michael Heinrich, Jörg Schmiedmayer and Raju Venugopalan. L.B., J.B. and S.F. acknowledge support by the DFG under the Collaborative Research Center SFB 1225 ISOQUANT (Project-ID 27381115) and the Heidelberg STRUCTURES Excellence Cluster under Germany’s Excellence Strategy EXC2181/1-390900948.

### v2

@@ VII Conclusions
In this work we studied the dynamics and particle production from a string between a quark-antiquark pair produced, for example, in an electron-positron collision. As a model, we used the bosonized version of the massive Schwinger model, which makes dynamics particularly easy to simulate numerically. We examine the dynamics of the coherent field and the particle production. Remarkably, the total particle number per rapidity interval to the fermion mass for highly massive or weakly coupled fermions turns out to be well represented by a Boltzmann factor.
@@
Very interesting extensions of this work include the consideration of flavor and color numbers, the incorporation of the production description beyond mesons, and taking into account baryons. Moreover, for the bounce scenario discussed in Sec. IV, the precise determination of the quantum state of fluctuations after the bounce would be highly valuable. The question of whether this can be achieved through analytic continuation from a calculation in Euclidean space is left for future work.
@@ ACKNOWLEDGEMENTS
For discussions on related topics, the authors would like to thank Sebastian Erne, Michael Heinrich, Jörg Schmiedmayer, and Raju Venugopalan. L.B., J.B., and S.F. acknowledge support by the DFG under the Collaborative Research Center SFB 1225 ISOQUANT (Project-ID 27381115) and the Heidelberg STRUCTURES Excellence Cluster under Germany’s Excellence Strategy EXC2181/1-390900948.
@@
APPENDIX B Renormalization
This appendix completes Sec. II, specifically providing explanation based on the functional renormalization group method to justify the renormalized potential (Eq. (5)) The theory in Eq. (1) is alternatively described by the microscopic action
with the scalar potential
$u_{\Lambda}=\frac{\exp(\gamma)}{2\pi}\Lambda m,$ (46)
The bosonized theory in Eq. (44) is an interacting quantum field theory subject to renormalization. A detailed discussion of renormalization in the functional formulation, addressing states with zero and nonzero temperature, can be found in Ref. [22]. The path integral systematically includes quantum fluctuations into correlation functions when moving from UV to IR, where the scale-dependent fluctuations enter corresponding coupling constants of the scale-dependent effective action $\Gamma_{k}$. In the limit $k\gg\Lambda$ all quantum fluctuations are suppressed, such that $\Gamma_{k}[\phi]=S_{\Lambda}[\phi]$. On the other side, for $k\to 0$, the flowing action would approach the full quantum effective action $\Gamma_{k=0}[\phi]=\Gamma[\phi]$, see Ref. [22] for a more detailed discussion. The regulator function is chosen such that $\Gamma_{k}$ smoothly interpolates between the microscopic action $S_{\Lambda}$ for $k=\Lambda$ and the macroscopic action $\Gamma$ for $k=0$. We assume that $\Gamma_{k}[\phi]$ is of the same form as Eq. (44) with potential term as in Eq. (45), except for the replacement $u_{\Lambda}\to u_{k}$, whereas the derivative term keeps its bare form. For the purpose of the present work, we shall be satisfied with this simplified approach where only a single parameter, the coupling $u_{\Lambda}$, gets replaced by the renormalized coupling $u$. There is, in fact, a good reason why $M$ does not need to be renormalized, which is explained below. The $k$-dependent effective potential $V_{k}(\phi)$ satisfies the exact equation [77],
$\partial_{k}V_{k}(\phi)=$ $\frac{1}{2}\int\limits_{p^{2}<\Lambda^{2}}\frac{\mathrm{d}^{2}p}{(2\pi)^{2}}\frac{2k}{p^{2}+k^{2}+V_{k}^{(2)}(\phi)}$ (47)
$=$ $\frac{k}{4\pi}\ln\left(\frac{\Lambda^{2}+k^{2}+V_{k}^{(2)}(\phi)}{k^{2}+V_{k}^{(2)}(\phi)}\right),$
with
To obtain a flow equation for $u_{k}$, we use the inverse Fourier expansion scheme,
This can be integrated with respect to $k$ to yield
$u=u_{0}=u_{\Lambda}\frac{M}{\Lambda}.$ (52)
The final form of the renormalized potential $V(\phi)$ is given by Eq. (5).